Twice the greater of two consecutive odd numbers is 13 less than three times the lesser. Find the integers.
Can someone please help. i did the problem and got -9 and -7 but im not sure.

Found 2 solutions by algebrahouse.com, jorel1380:
Answer by algebrahouse.com(1659)

"Twice the greater of two consecutive odd numbers is 13 less than three times the lesser. Find the integers."

x = 1st odd integer
x + 2 = 2nd odd integer {odd numbers increase by 2 each time}

2(x + 2) = 3x - 13 {twice the greater is 13 less than three times the lesser}
2x + 4 = 3x - 13 {used distributive property}
4 = x - 13 {subtracted 2x from both sides}
x = 17 {added 13 to both sides}
x + 2 = 19 {substituted 17, in for x, into x + 2}

17 and 19 are the two odd integers

For more help from me, visit: www.algebrahouse.com



Answer by jorel1380(3719)

3m=2(m+2)+13
3m=2m+17
m=17
m+2=19
The two numbers are 17 and 19..